# DeCode: Decoupling Content and Delivery for Medical QA

### v2

@@ Abstract
Large language models (LLMs) exhibit strong medical knowledge and can generate factually accurate responses. However, existing models often fail to account for individual patient contexts, producing answers that are clinically correct yet poorly aligned with patients’ needs. In this work, we introduce DeCode (Decoupling Content and Delivery), a training-free, model-agnostic framework that adapts existing LLMs to produce contextualized answers in clinical settings. We evaluate DeCode on OpenAI HealthBench, a comprehensive and challenging benchmark designed to assess clinical relevance and validity of LLM responses. DeCode improves the previous state-of-the-art from $28.4\%$ to $49.8\%$, corresponding to a $75\%$ relative improvement. Experimental results suggest the effectiveness of DeCode in improving clinical question answering of LLMs.
@@ 1 Introduction
Importantly, DeCode operates in a training-free paradigm, orchestrating the generation process through explicit clinical formulation and structured discourse constraints. Empirically, we demonstrate that DeCode substantially improves performance on HealthBench, increasing the prior state-of-the-art score from $28.4\%$ to $49.8\%$. Furthermore, we show that DeCode generalizes consistently across multiple leading LLMs, suggesting that the framework captures model-agnostic principles for personalized medical response generation.
@@ 4.2 Implementation Details
We compare against: (i) Zero-shot, which directly prompts the base LLM using the conversation history; (ii) MDAgents (Kim et al., 2024), which recruits specialized experts based on query complexity and aggregates their responses; (iii) KAMAC (Wu et al., 2025b), which dynamically recruits experts during generation; and (iv) MuSeR (Zhou et al., 2025), which applies a self-refinement generation strategy. For KAMAC, the model selects the initial number of experts and the discussion is fixed to two rounds. For MuSeR, we use the authors’ self-refinement prompts at inference time, without model distillation or fine-tuning. A detailed cost and latency breakdown for all frameworks is included in the appendix.
@@ 5.3 Comparison with LLM Frameworks
Table 3: Comparison with leading LLM-based frameworks. We evaluate MDAgents Kim et al. (2024), KAMAC Wu et al. (2025b), and MuSeR Zhou et al. (2025), together with standard zero-shot prompting and DeCode, all using OpenAI o3 on the HealthBench hard subset. Inline deltas are reported relative to the zero-shot baseline, with improvements greater than 2 points highlighted.
Zero-Shot MDAgents KAMAC MuSeR DeCode
Metric ($\uparrow$, %) (Kim et al., 2024) (Wu et al., 2025b) (Zhou et al., 2025) (Ours)
@@
Emergency Referrals 27.0 36.3 (+9.3) 33.1 (+6.1) 53.3 (+26.3) 59.1 (+32.1)
Context Seeking 30.0 34.4 (+4.4) 27.4 (-2.6) 60.1 (+30.1) 58.3 (+28.3)
Global Health 31.8 43.0 (+11.2) 30.2 (-1.6) 44.4 (+12.6) 49.8 (+18.0)
Health Data Tasks 17.0 19.1 (+2.1) 7.7 (-9.3) 37.2 (+20.2) 35.6 (+18.6)
Communications 29.2 38.0 (+8.8) 32.9 (+3.7) 40.4 (+11.2) 43.8 (+14.6)
Hedging 30.9 39.2 (+8.3) 29.9 (-1.0) 51.2 (+20.3) 54.6 (+23.7)
Complex Responses 24.4 31.6 (+7.2) 28.5 (+4.1) 37.7 (+13.3) 41.3 (+16.9)
@@
Accuracy 45.6 52.9 (+7.3) 43.8 (-1.8) 53.5 (+7.9) 54.3 (+8.7)
Completeness 30.7 45.8 (+15.1) 36.1 (+5.4) 49.1 (+18.4) 58.8 (+28.1)
Communication Quality 55.5 49.5 (-6.0) 46.7 (-8.8) 54.1 (-1.4) 54.2 (-1.3)
Context Awareness 4.0 4.6 (+0.6) 0.0 (-4.0) 46.0 (+42.0) 40.5 (+36.5)
Instruction Following 45.8 53.5 (+7.7) 36.2 (-9.6) 46.2 (+0.4) 46.5 (+0.7)
In this experiment, we evaluate representative LLM-based medical QA frameworks on the HealthBench hard subset using OpenAI o3 as the base LLM. Specifically, we consider MDAgents Kim et al. (2024), KAMAC Wu et al. (2025b), and MuSeR Zhou et al. (2025) with results summarized in Table 3.
Relative to the zero-shot baseline, MDAgents demonstrates consistent improvements across all conversation themes and four of the five evaluation axes. Notably, it achieves the strongest performance on instruction following among all compared methods. This behavior can be attributed to its complexity-driven orchestration strategy: MDAgents first estimates the difficulty of a given medical query and determines whether it can be handled by a single agent or requires a coordinated team of specialized experts. Once the team composition is selected, it remains fixed throughout the generation process. This upfront complexity assessment enables stable role assignment and coherent multi-agent collaboration, which appears particularly effective for instruction-heavy medical QA scenarios.
In contrast, KAMAC does not yield consistent gains over the zero-shot baseline and, in several cases, exhibits notable degradations in context awareness, communication quality, and instruction following. Unlike MDAgents, KAMAC follows a knowledge-driven strategy that dynamically recruits new specialists during the generation process when existing agents identify missing domain knowledge. While this adaptive recruitment mechanism is intended to enhance coverage, our analysis suggests that introducing new experts mid-stream can disrupt conversational coherence. Specifically, the newly added agents often generate responses that overlap with existing contributions or shift the discussion focus, leading to redundancy and task-level confusion. These effects are amplified in longer, multi-round discussions, ultimately degrading response quality.
@@
Taken together, these results indicate that multi-agent medicalQA performance depends critically on how contextual information and communicative intent are structured during generation. While fixed-team orchestration and self-refinement improve context sensitivity, dynamically introducing new experts can incur coordination overhead that degrades coherence. Explicitly separating clinical content from discourse strategy provides a more stable alternative, enabling balanced optimization of medical accuracy and communication quality, as demonstrated by DeCode.
@@ Appendix A Cost & Latency Analysis
In this section, we analyze the cost and latency of the DeCode framework. In addition, we include the cost and latency of MDAgents Kim et al. (2024), KAMAC Wu et al. (2025b), and MuSeR Zhou et al. (2025). To ensure fair comparison, we assume an output token rate of $70$ tokens/second to avoid latency fluctuations due to API re-routing. The details of the DeCode framework is listed in Table  5. Comparisons between the various LLM-based medical QA frameworks are provided in Table 6.
@@ Appendix B Prompt Templates
- Economic condition (inferred from context)

### v3

@@ Abstract
Large language models (LLMs) exhibit strong medical knowledge and can generate factually accurate responses. However, existing models often fail to account for individual patient contexts, producing answers that are clinically correct yet poorly aligned with patients’ needs. In this work, we introduce DeCode (Decoupling Content and Delivery), a training-free, model-agnostic framework that adapts existing LLMs to produce contextualized answers in clinical settings. We evaluate DeCode on OpenAI HealthBench, a comprehensive and challenging benchmark designed to assess clinical relevance and validity of LLM responses. DeCode boosts zero-shot performance from $28.4\%$ to $49.8\%$ and achieves new state-of-the-art compared to existing methods. Experimental results suggest the effectiveness of DeCode in improving clinical question answering of LLMs.
@@ 1 Introduction
Importantly, DeCode operates in a training-free paradigm, orchestrating the generation process through explicit clinical formulation and structured discourse constraints. Empirically, we demonstrate that DeCode substantially improves performance on HealthBench, boosting zero-shot performance from $28.4\%$ to $49.8\%$ and surpassing the recent state-of-the-art, MuSeR ($47.1\%$), by an absolute margin of $2.7\%$. Furthermore, we show that DeCode generalizes consistently across multiple leading LLMs, suggesting that the framework captures model-agnostic principles for personalized medical response generation.
@@ 4.2 Implementation Details
We compare against: (i) standard Zero-shot (ii) Single Stage, an improved baseline that compresses the four phases of DeCode into a single prompt; (iii) MDAgents (Kim et al., 2024), which recruits specialized experts based on query complexity and aggregates their responses; (iv) KAMAC (Wu et al., 2025b), which dynamically recruits experts during generation; and (v) MuSeR (Zhou et al., 2025), which applies a self-refinement generation strategy. For MuSeR, we use the authors’ self-refinement prompts at inference time, without model distillation or fine-tuning. A detailed cost and latency breakdown for all frameworks is included in the appendix.
@@ 5.3 Comparison with LLM Frameworks
Table 3: Comparison with leading LLM-based frameworks. We evaluate MDAgents, KAMAC, and MuSeR, together with standard zero-shot prompting, a single-stage baseline, and DeCode. All methods use OpenAI o3 on the HealthBench hard subset. Inline deltas are reported relative to the zero-shot baseline, with improvements greater than 2 points highlighted.
Metric ($\uparrow$, %) Zero-Shot Single Stage MDAgents KAMAC MuSeR DeCode
@@
Emergency Referrals 27.0 53.4 (+26.4) 36.3 (+9.3) 33.1 (+6.1) 53.3 (+26.3) 59.1 (+32.1)
Context Seeking 30.0 44.8 (+14.8) 34.4 (+4.4) 27.4 (-2.6) 60.1 (+30.1) 58.3 (+28.3)
Global Health 31.8 38.7 (+6.9) 43.0 (+11.2) 30.2 (-1.6) 44.4 (+12.6) 49.8 (+18.0)
Health Data Tasks 17.0 24.0 (+7.0) 19.1 (+2.1) 7.7 (-9.3) 37.2 (+20.2) 35.6 (+18.6)
Communications 29.2 33.6 (+4.4) 38.0 (+8.8) 32.9 (+3.7) 40.4 (+11.2) 43.8 (+14.6)
Hedging 30.9 45.2 (+14.3) 39.2 (+8.3) 29.9 (-1.0) 51.2 (+20.3) 54.6 (+23.7)
Complex Responses 24.4 33.3 (+8.9) 31.6 (+7.2) 28.5 (+4.1) 37.7 (+13.3) 41.3 (+16.9)
@@
Accuracy 45.6 43.6 (-2.0) 52.9 (+7.3) 43.8 (-1.8) 53.5 (+7.9) 54.3 (+8.7)
Completeness 30.7 44.5 (+13.8) 45.8 (+15.1) 36.1 (+5.4) 49.1 (+18.4) 58.8 (+28.1)
Communication Quality 55.5 59.9 (+4.4) 49.5 (-6.0) 46.7 (-8.8) 54.1 (-1.4) 54.2 (-1.3)
Context Awareness 4.0 29.9 (+25.9) 4.6 (+0.6) 0.0 (-4.0) 46.0 (+42.0) 40.5 (+36.5)
Instruction Following 45.8 44.7 (-1.1) 53.5 (+7.7) 36.2 (-9.6) 46.2 (+0.4) 46.5 (+0.7)
In this experiment, we evaluate representative LLM-based medical QA frameworks on the HealthBench hard subset using OpenAI o3 as the base LLM. Specifically, we consider a Single Stage baseline, MDAgents Kim et al. (2024), KAMAC Wu et al. (2025b), and MuSeR Zhou et al. (2025), with results summarized in Table 3.
To isolate DeCode’s structural benefits, we introduce a single-stage baseline that compresses our framework into a single prompt, forcing the LLM to internally process all steps before responding. Although the single-stage baseline improves over zero-shot prompting, it still lags behind DeCode by a significant margin. This gap highlights the importance of decoupling modules: requiring a single LLM call to simultaneously manage context extraction, clinical reasoning, and discourse planning exceeds what the model can reliably coordinate within one generation pass. By explicitly separating these tasks, DeCode systematically reduces this burden, achieving the highest overall performance among all evaluated methods.
Relative to the zero-shot baseline, MDAgents demonstrates consistent improvements across all conversation themes and four of the five evaluation axes. Notably, it achieves the strongest performance on instruction following among all compared methods. This success stems from its complexity-driven orchestration strategy: by estimating query difficulty upfront to form a fixed expert team, it ensures stable role assignment and coherent multi-agent collaboration. This static team composition appears particularly effective for instruction-heavy medical QA scenarios.
In contrast, KAMAC yields inconsistent gains over the zero-shot baseline and exhibits notable degradations in context awareness, communication quality, and instruction following. Its knowledge-driven strategy dynamically recruits specialists mid-generation to address missing domain knowledge. While this adaptive recruitment mechanism is intended to enhance coverage, our analysis suggests that introducing new experts mid-stream can disrupt conversational coherence. Specifically, the newly added agents often generate responses that overlap with existing contributions or shift the discussion focus, leading to redundancy and task-level confusion. These effects are amplified in longer, multi-round discussions, ultimately degrading response quality.
@@
Taken together, these results indicate that medical QA performance depends critically on how contextual information and communicative intent are structured during generation. While fixed-team orchestration and self-refinement improve context sensitivity, dynamic expert recruitment incurs overhead. Explicitly decoupling clinical content from discourse strategy provides a more stable architecture, enabling balanced optimization of medical accuracy and communication quality, as demonstrated by DeCode.
@@ Appendix A Cost & Latency Analysis
In this section, we analyze the cost and latency of the DeCode framework. In addition, we include the cost and latency of MDAgents Kim et al. (2024), KAMAC Wu et al. (2025b), and MuSeR Zhou et al. (2025). To ensure fair comparison, we assume an output token rate of $70$ tokens/second to avoid latency fluctuations due to API re-routing. Furthermore, costs are estimated based on the official OpenAI o3 pricing model, $\mathdollar 2.00$ per 1M input tokens and $\mathdollar 8.00$ per 1M output tokens.
DeCode Component Breakdown.
Table 5 details the token usage and latency of each module within our framework. By executing the Profiler (Background and User Needs) and the Formulator modules concurrently, we effectively mask the latency of the initial extraction steps. The final Synthesizer accounts for the largest portion of the overall latency ($20.83$ s), as it is responsible for generating the comprehensive, patient-facing response. Overall, the parallelized architecture strictly bounds the total latency to approximately $45$ seconds per query.
Comparative Efficiency.
Table 6 highlights the substantial computational overhead commonly observed in complex multi-agent systems. Frameworks like KAMAC and MDAgents incur massive token overheads due to redundant context-sharing and lengthy debate loops. This results in prohibitively high costs and impractical latencies. In contrast, DeCode’s decoupled, linear pipeline avoids the extensive communication overhead inherent in multi-agent frameworks. Consequently, DeCode achieves the lowest average cost ($\mathdollar 0.037$ per query) and latency ($45.47$ s) among all evaluated frameworks.
@@ Appendix B Prompt Templates
- Financial Constraints (inferred from context)
@@ 2. WHAT NOT TO DO/COVER (NOT TO DO):
Appendix C Case Study Illustration
To provide a qualitative assessment of DeCode, we present a representative case directly from the HealthBench dataset.
[…]
“feeling depressed for the last few months”
Baseline: Zero-Shot (o3) | Score: 0.259
[…]
1. Talk to a professional: [Generic advice omitted]
2. Reach out for support: [Generic advice omitted]
3. Everyday self-care strategies: [Generic advice omitted]
4. Keep track of feelings: [Generic advice omitted]
5. Crisis resources: If you ever feel like you might act on thoughts of harming yourself, call emergency services….
Ours: DeCode (o3) | Score: 0.907
[…]
• Have you had any thoughts of ending your life or harming yourself?
If you feel you might act on these thoughts, please call 911 right now…
To understand what you’re experiencing and guide next steps, could you tell me a little more?
1. Symptoms & impact: When did the low mood start? How is it affecting work, school, or self-care?
2. Safety & red flags: Any thoughts of harming others? Have you noticed hearing or seeing things others don’t?
3. Bipolar screening: Have you ever had periods of unusually high energy (mania)?
Practical Next Steps: Schedule a visit with a primary-care doctor. Evidence-based talk therapies (CBT) and medications may be appropriate. Keep a regular routine and stay connected.
Table 7: Qualitative comparison on a mental health query. DeCode (blue) prioritizes emergency triage, whereas the baseline (red) provides passive advice.
Qualitative Comparison.
According to the HealthBench grading results, the zero-shot baseline incurred a severe penalty (10-10 points) for failing to assess immediate suicide risk, defaulting to passive advice. In contrast, DeCode earned full points (+10+10 points) for context awareness by explicitly screening for suicidal ideation and red flags. While both models received a verbosity penalty (5-5 points) for providing over six suggestions, DeCode intentionally trades conversational brevity for comprehensive clinical safety and proper triage, yielding a vastly superior overall score ($0.907$ vs. $0.259$).